what would and equation be to the math problem where there is a rectangle with a perimeter of 84 feet and the width is 8 feet less than the length?
Answer by mestrydilip(12)

Let Length be x feet
Width = x - 8
Perimeter = 2(Length + Width)
84 = 2(x + x - 8)
42 = 2x - 8
50 = 2x
x = 25
Length = 25 feet
Width = 25 - 8 = 17 feet